Andrea Deposits $1000 into a savings account today. The annual interest rate is 10%, but compounds semi-annually. How many years will it take for her money to grow to $4,356,561?

A. 171.7 years

B. 87.9 years

C. 85.9 years

D. 175.8 years
Answer #1

Semi annual Rate =10%/2 =5%
PV =1000
FV =4356561
Number of Years using excel formula =NPER(5%,0,-1000,4356561)/2 =85.9 years

alternative method using financial calculator
I/Y =5%, PV =-1000;FV =4356561; CPT N =171.74
Number of Years =171.74/2 =85.9 years

Option c is correct option